# Measurability in the Fundamental Theorem of Statistical Learning

### v2

@@ 4 Connections to Model Theory
To specify a (first-order) language $\mathcal{L}$, we simply list the collection of function, relation and constant symbols. For instance, $\mathcal{L}_{\mathrm{r}}=\{+,-,\cdot,0,1\}$ denotes the language of rings, where +,,+,-$\cdot$+ , - , ⋅ are binary function symbols and $0,1$ are constant symbols, $\mathcal{L}_{<}=\{<\}$ denotes the language of orderings, where $<$ is a binary relation symbol, the language of ordered rings is given by $\mathcal{L}_{\mathrm{or}}=\mathcal{L}_{\mathrm{r}}\cup\mathcal{L}_{<}$, and the language of ordered exponential fields is given by $\mathcal{L}_{\exp}=\mathcal{L}_{\mathrm{or}}\cup\{\exp\}$ with $\exp$ denoting a unary function symbol. For a language $\mathcal{L}$, an $\mathcal{L}$–structure $\mathcal{M}$ is given by a non-empty domain $M$ and an interpretation of each symbol in $\mathcal{L}$. For instance, $\mathbb{R}_{\mathrm{or}}=(\mathbb{R},+,\cdot,-,0,1,<)$ denotes the ordered field of real numbers (with standard interpretation of the symbols), and its expansion $\mathbb{R}_{\exp}=(\mathbb{R},+,\cdot,-,0,1,<,\exp)$ denotes the real exponential field with standard exponential $\exp\colon x\mapsto\mathrm{e}^{x}$.
@@
$(x<0\to y=0)\wedge(0\leq x\to y=x).\vspace{-0.7em}$
Given a language $\mathcal{L}$ expanding $\mathcal{L}_{<}$, and an $\mathcal{L}$-structure $\mathcal{M}$ whose reduct $(M,<)$ is a linear order, recall that $\mathcal{M}$ is o-minimal if any set $A\subseteq M$ that is definable over $\mathcal{M}$ can be expressed as a finite union of points and open intervals (see Marker [20, Definition 3.1.18]). Steinhorn [31, page 27] notes that “PAC learning is directly relevant to neural networks whose architecture defines concept classes that consist of definable families of definable sets in o-minimal structures.” Most crucially, due to Wilkie [35] the real exponential field $\mathbb{R}_{\exp}$ is an o-minimal $\mathcal{L}_{\exp}$–structure. O-minimality implies the property NIP (see Pillay and Steinhorn [23, Corollary 3.10]), which in turn ensures that hypothesis spaces of the form $\mathcal{H}_{\mathcal{X}}^{\varphi}$ have finite VC dimension. All hypothesis spaces as in Example 4.2 are therefore of finite VC dimension. As mentioned in the introduction, these observations were first made by Laskowski [18]. Since Laskowski’s original arguments dealt with quantitative bounds for concept classes rather than hypothesis spaces over definable instance spaces $\mathcal{X}$, we present here a self-contained and more direct proof of these observations. We first introduce the relevant model-theoretic property NIP (see Poizat [24, § 12.4]), which is originally due to Shelah [28].
@@
⊧̸$\mathcal{M}\not\models\bigwedge_{i\in J}\varphi(\underline{a}_{i};\underline{w}_{J})\wedge\bigwedge_{i\in[m]\setminus J}\neg\varphi(\underline{a}_{i};\underline{w}_{J}).$
@@
Definition 4.3 expresses that a formula $\varphi(\underline{x};\underline{p})$ has NIP if there is $m\in\mathbb{N}$ such that no set of size $m$ is shattered by $\varphi$ (see Simon [30, Remark 2.3]).
@@
Proposition 4.5 closely relates the model-theoretic notion NIP with the concept of VC dimension from Statistical Learning Theory. Namely, any hypothesis space $\mathcal{H}^{\varphi}_{\mathcal{X}}$ over an NIP structure has finite VC dimension. On the other hand, the Fundamental Theorem of Statistical Learning connects VC dimension with PAC learning. Therefore, drawing a direct connection between NIP and PAC learning suggests itself (cf. e.g. Steinhorn [31, page 27]). However, the measurability requirements that we carefully scrutinized in Section 3 have to be accounted for.
In the following, we examine when hypothesis spaces $\mathcal{H}^{\varphi}_{\mathcal{X}}$ are PAC learnable for the special case that the underlying $\mathcal{L}$–structure is an o-minimal expansion of $\mathbb{R}_{\mathrm{or}}=(\mathbb{R},+,\cdot,-,0,1,<)$. While our measure-theoretic examination is guided by the work of Karpinski and Macintyre [17], we add further adjustments to establish agnostic PAC learnability in this context.
@@
$f_{U}\colon\mathcal{Z}^{m}\times\mathbb{R}^{\ell}\to[0,1],\ (\underline{z};\underline{w})\mapsto\big{|}\operatorname{er}_{\mathbb{D}}(h_{\underline{w}})-\hat{\operatorname{er}}_{\underline{z}}(h_{\underline{w}})\big{|}$
@@
Thus, the map $\hat{\operatorname{er}}_{\square}(h_{\square})$ is $(\Sigma_{\mathcal{Z}}^{m}\otimes\mathcal{B}(\mathbb{R}^{\ell}))$–measurable. The measurability of the maps $\operatorname{er}_{\mathbb{D}}(h_{\square})$ and $\hat{\operatorname{er}}_{\square}(_{\square})$ implies that the maps $f_{U}$ and $f_{V}$ are measurable. Since we assume that the probability space $(\mathcal{Z}^{k},\Sigma_{\mathcal{Z}}^{k},\mathbb{D}^{k})$ is complete for any $k\in\mathbb{N}$, we can apply the arguments of Pollard [25, page 197] to deduce that the map
@@
    Given an arbitrary probability space $(\Omega,\Sigma,\mathbb{P})$, there always exists a (Lebesgue) completion, i.e. a probability space $(\Omega,\hat{\Sigma},\hat{\mathbb{P}})$ that is complete and fulfills $\Sigma\subseteq\hat{\Sigma}$ as well as ^$\hat{\mathbb{P}}\mathord{\upharpoonright}_{\Sigma}=\mathbb{P}$ (cf. Bogachev [6, page 22]). However, it is well-known that the product of two complete probability spaces is in general not complete. Therefore, considering the completion $(\mathcal{Z},\hat{\Sigma}_{\mathcal{Z}},\hat{\mathbb{D}})$ does not guarantee that also the product space $(\mathcal{Z}^{m},(\hat{\Sigma}_{\mathcal{Z}})^{m},(\hat{\mathbb{D}})^{m})$ is complete for any $m\in\mathbb{N}$.
@@
    Applying similar arguments to demonstrate the measurability of the map $U=U(\mathcal{H}_{\mathcal{X}}^{\varphi},m,\mathbb{D})$ would require dealing with sets of the form
    $\{\underline{w}\in\mathbb{R}^{\ell}\mid\operatorname{er}_{\mathbb{D}}(h_{\underline{w}})\geq\alpha\}=\{\underline{w}\in\mathbb{R}^{\ell}\mid\mathbb{D}(\mathcal{Z}\setminus\Gamma(h_{\underline{w}}))\geq\alpha\}$
    with $\alpha\in(0,1)$, which are a priori not $\mathcal{L}$–definable.
@@
    Theorem 4.7 can potentially be extended to other structures that are, in some sense, tame (e.g. o-minimal structures, structures with quantifier elimination) and naturally endowed with a topology (cf. [17, § 5]). However, one must carefully examine whether the arguments of Pollard [25, page 197] – or those upon which Pollard’s approach relies (cf. Dellacherie and Meyer [9, Chapter III]) – can still be applied in such settings. These arguments, for instance, require the topological space to be metrizable, separable and locally compact.
@@ 5 Further Work
Appendix A Appendix
@@ A.1 Auxiliary Lemmas
is a random variable on the discrete measurable space ({±$(\{\pm 1\}^{m},\mathcal{P}(\{\pm 1\}^{m}))$ with
@@
Furthermore, $V_{h}$ is an average of independent random variables, each of which takes values in $[-1,1]$.
@@
where, for any $i\in[m]$, the random variable [^12] Note that the map $V_{h}^{(i)}$ is measurable, as it is defined on the discrete measurable space ({±$(\{\pm 1\}^{m},\mathcal{P}(\{\pm 1\}^{m}))$. $V_{h}^{(i)}$ is given by
@@
$\mathbb{P}(V_{h}^{(i)}=0)$ =$=\begin{cases}1&\text{if }\Delta_{h}^{(i)}=0,\\ 0&\text{otherwise},\end{cases}$
$\mathbb{P}(V_{h}^{(i)}=\ell)$ =$=\begin{cases}\frac{1}{2}&\text{if }\Delta_{h}^{(i)}\neq 0,\\ 0&\text{otherwise},\end{cases}\quad\quad(\ell\in\{\pm 1\}),$
@@
The following result is a refined version of Shalev-Shwartz and Ben-David [27, Lemma A.4].
@@
If $\beta\geq e$, then applying [27, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=\beta$) yields
@@
Otherwise, applying [27, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=e$) yields
@@
Let $(\Omega_{1},\Sigma_{1})$,$(\Omega_{2},\Sigma_{2})$ be measurable spaces and let $A_{1}\subseteq\Omega_{1}$, $A_{2}\subseteq\Omega_{2}$. Then we have $(\Sigma_{1}\otimes\Sigma_{2})\cap(A_{1}\times A_{2})=(\Sigma_{1}\cap A_{1})\otimes(\Sigma_{2}\cap A_{2})$.
@@
=\;= $\{(E_{1}\cap A_{1})\times(E_{2}\cap A_{2})\mid E_{1}\in\Sigma_{1},E_{2}\in\Sigma_{2}\}$
=\;= $\{E^{\prime}_{1}\times E^{\prime}_{2}\mid E^{\prime}_{1}\in(\Sigma_{1}\cap A_{1}),E^{\prime}_{2}\in(\Sigma_{2}\cap A_{2})\},$
@@ A.2 Well-Behaved Hypothesis Spaces
$\underline{z}$ $\mapsto\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}$
@@
is $\Sigma_{\mathcal{Z}}^{2m}$–measurable.
In general, we however need to take suprema over uncountable families of these measurable functions. In this situation, verifying the measurability of the maps $U$ and $V$ is more complicated or even impossible (see Example A.13). However, in this section we aim at identifying further sufficient conditions that might be helpful in the uncountable setting as well.
We now introduce the concept of a universally separable hypothesis space. Dudley [11, page 902] was the first to introduce this property “as a way of avoiding measurability difficulties” (Pollard [25, page 38]).
@@
We next prove that universal separability ensures measurability of the maps $U$ and $V$ from Definition 3.2. Similar statements in a more general setting can be found in the literature (cf. e.g. Ben-David, Mansour and Benedek [3, Lemma 5.2] and Pollard [25, Chapter II, Problem 3]). Guided by the proof of [3, Lemma 5.2] and by arguments suggested by D. Pollard [^13] The second author thanks D. Pollard for these suggestions via private communication., we present here a detailed proof adapted to the maps $U$ and $V$.
@@
[…]
@@
$\big{|}\operatorname{er}_{\mathbb{D}}(h_{n})-\operatorname{er}_{\mathbb{D}}(h)\big{|}=\big{|}\mathbb{D}(\Gamma(h))-\mathbb{D}(\Gamma(h_{n}))\big{|}\leq\mathbb{D}(\Gamma(h)\triangle\Gamma(h_{n})),$
@@
The universal separability condition applies to many standard examples of hypothesis spaces (cf. Pollard [25, Chapter II, Problems 4, 5 and 7]). However, universal separability is no necessary condition for ensuring well-behavedness. In fact, Example A.15 presents a quite simple hypothesis space that is not universally separable, but well-behaved with respect to certain sets of distributions.
Further considerations on the measurability of suprema over uncountable families of measurable functions can be found in Dudley [11], [12, Chapter 10], Gaenssler [13, Chapter 4] and Pollard [25, Appendix C]. For instance, [25, Appendix C] introduces the concept of permissible function classes, which Haussler [15, Section 9.2] also takes up. Moreover, recall that the arguments in [25, Appendix C] constitute the core for our proof of Theorem 4.7.
@@
    $C\subseteq\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}|\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)|\leq\varepsilon\right.\right\}$
    and $\text{and }\mathbb{D}^{m}(C)\geq 1-\delta.$
@@
    For the proofs of Theorem 3.6 and, building on it, Theorem 3.8 both $U$ and $V$ have to measurable, as a key step is to bound their expected values. To adjust these proofs for arbitrary hypothesis spaces, we first note that the maps $U$ and $V$ are certainly $\mathcal{P}(\mathcal{Z}^{m})$–measurable and $\mathcal{P}(\mathcal{Z}^{2m})$–measurable, respectively. If $\mathbb{D}$ is discrete, then we can write
    $\mathbb{D}=\sum\limits_{z\in C_{\mathbb{D}}}p_{z}\delta_{z}$
    for suitable $p_{z}\in[0,1]$ and a finite set $C_{\mathbb{D}}\subseteq\mathcal{Z}$, and we can extend [^14] It is easily verified that the restriction of $\hat{\mathbb{D}}$ to $\Sigma_{\mathcal{Z}}$ coincides with $\mathbb{D}$. $\mathbb{D}$ to a distribution $\hat{\mathbb{D}}$ on $\mathcal{P}(\mathcal{Z})$ by setting
    $\hat{\mathbb{D}}(C):=\mathbb{D}(\underbrace{C\cap C_{\mathbb{D}}}_{\in\Sigma_{\mathcal{Z}}})$
    for $C\subseteq\mathcal{Z}$. In our computations, we can then replace the probability space $(\mathcal{Z},\Sigma_{\mathcal{Z}},\mathbb{D})$ and its powers $(\mathcal{Z}^{k},\Sigma_{\mathcal{Z}}^{k},\mathbb{D}^{k})$ by the “discretized” probability space $(\mathcal{Z},\mathcal{P}(\mathcal{Z}),\hat{\mathbb{D}})$ and its powers $(\mathcal{Z}^{k},\mathcal{P}(\mathcal{Z}^{k}),\hat{\mathbb{D}}^{k})$, in order to omit measurability difficulties. With these replacements, following the proofs of Theorem 3.6 and Theorem 3.8, we then obtain
@@
    for the finite set $C=\{\underline{z}\in C_{\mathbb{D}}^{m}\mid U(\underline{z})\leq\varepsilon(m,\delta)\}\in\Sigma_{\mathcal{Z}}$ and any $m\geq m_{0}$, where
    $m_{0}=m_{0}(\varepsilon,\delta)=\lceil\max\{m_{0}^{(1)},m_{0}^{(2)},m_{0}^{(3)}\}\rceil.$
@@
    $\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\frac{\varepsilon}{4}\right.\right\}$

### v3

@@ 4 Connections to Model Theory
To specify a (first-order) language $\mathcal{L}$, we simply list the collection of function, relation and constant symbols. For instance, $\mathcal{L}_{\mathrm{r}}=\{+,-,\cdot,0,1\}$ denotes the language of rings, where $+,-,\cdot$ are binary function symbols and $0,1$ are constant symbols, $\mathcal{L}_{<}=\{<\}$ denotes the language of orderings, where << is a binary relation symbol, the language of ordered rings is given by $\mathcal{L}_{\mathrm{or}}=\mathcal{L}_{\mathrm{r}}\cup\mathcal{L}_{<}$, and the language of ordered exponential fields is given by $\mathcal{L}_{\exp}=\mathcal{L}_{\mathrm{or}}\cup\{\exp\}$ with $\exp$ denoting a unary function symbol. For a language $\mathcal{L}$, an $\mathcal{L}$–structure $\mathcal{M}$ is given by a non-empty domain $M$ and an interpretation of each symbol in $\mathcal{L}$. For instance, $\mathbb{R}_{\mathrm{or}}=(\mathbb{R},+,-,\cdot,0,1,<)$ denotes the ordered field of real numbers (with standard interpretation of the symbols), and its expansion $\mathbb{R}_{\exp}=(\mathbb{R},+,-,\cdot,0,1,<,\exp)$ denotes the real exponential field with standard exponential $\exp\colon x\mapsto\mathrm{e}^{x}$.
@@
$(x<0\to y=0)\wedge(0\leq x\to y=x).\vskip-6.99997pt$
Given a language $\mathcal{L}$ expanding $\mathcal{L}_{<}$, and an $\mathcal{L}$-structure $\mathcal{M}$ whose reduct $(M,<)$ is a linear order, recall that $\mathcal{M}$ is o-minimal if any set $A\subseteq M$ that is definable over $\mathcal{M}$ can be expressed as a finite union of points and open intervals (see Marker [21, Definition 3.1.18]). Steinhorn [32, page 27] notes that “PAC learning is directly relevant to neural networks whose architecture defines concept classes that consist of definable families of definable sets in o-minimal structures.” Most crucially, due to Wilkie [37] the real exponential field $\mathbb{R}_{\exp}$ is an o-minimal $\mathcal{L}_{\exp}$–structure. O-minimality implies the property NIP (see Pillay and Steinhorn [24, Corollary 3.10]), which in turn ensures that hypothesis spaces of the form $\mathcal{H}_{\mathcal{X}}^{\varphi}$ have finite VC dimension. All hypothesis spaces as in Example 4.2 are therefore of finite VC dimension. As mentioned in the introduction, these observations were first made by Laskowski [18]. Since Laskowski’s original arguments dealt with quantitative bounds for concept classes rather than hypothesis spaces over definable instance spaces $\mathcal{X}$, we present here a self-contained and more direct proof of these observations. We first introduce the relevant model-theoretic property NIP (see Poizat [25, § 12.4]), which is originally due to Shelah [29].
@@
$\mathcal{M}\not\models\bigwedge_{i\in J}\varphi(\underline{a}_{i};\underline{w}_{J})\wedge\bigwedge_{i\in[m]\setminus J}\neg\varphi(\underline{a}_{i};\underline{w}_{J}).$
@@
Definition 4.3 expresses that a formula $\varphi(\underline{x};\underline{p})$ has NIP if there is $m\in\mathbb{N}$ such that no set of size $m$ is shattered by $\varphi$ (see Simon [31, Remark 2.3]).
@@
Proposition 4.5 closely relates the model-theoretic notion NIP with the concept of VC dimension from Statistical Learning Theory. Namely, any hypothesis space $\mathcal{H}^{\varphi}_{\mathcal{X}}$ over an NIP structure has finite VC dimension. On the other hand, the Fundamental Theorem of Statistical Learning connects VC dimension with PAC learning. Therefore, drawing a direct connection between NIP and PAC learning suggests itself (cf. e.g. Steinhorn [32, page 27]). However, the measurability requirements that we carefully scrutinized in Section 3 have to be accounted for.
In the following, we examine when hypothesis spaces $\mathcal{H}^{\varphi}_{\mathcal{X}}$ are PAC learnable for the special case that the underlying $\mathcal{L}$–structure is an o-minimal expansion of $\mathbb{R}_{\mathrm{or}}=(\mathbb{R},+,-,\cdot,0,1,<)$. While our measure-theoretic examination is guided by the work of Karpinski and Macintyre [17], we add further adjustments to establish agnostic PAC learnability in this context.
@@
$f_{U}\colon\mathcal{Z}^{m}\times\mathbb{R}^{\ell}\to[0,1],\ (\underline{z};\underline{w})\mapsto\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h_{\underline{w}})-\hat{\operatorname{er}}_{\underline{z}}(h_{\underline{w}})\big|$
@@
[…]
@@
[…]
@@
    Applying similar arguments to demonstrate the measurability of the map $U=U(\mathcal{H}_{\mathcal{X}}^{\varphi},m,\mathbb{D})$ would require dealing with sets of the form {wR^ℓ∣er_D(h_w)≥α}={wR^ℓ∣D(Z∖Γ(h_w))≥α} with $\alpha\in(0,1)$, which are a priori not $\mathcal{L}$–definable.
@@
    Theorem 4.7 can potentially be extended to other structures that are, in some sense, tame (e.g. o-minimal structures, structures with quantifier elimination) and naturally endowed with a topology (cf. [17, § 5]). However, one must carefully examine whether the arguments of Pollard [26, page 197] – or those upon which Pollard’s approach relies (cf. Dellacherie and Meyer [9, Chapter III]) – can still be applied in such settings. These arguments, for instance, require the topological space to be metrizable, separable and locally compact.
@@ 5 Further Work
Appendix A Appendix: Lemmas, Examples and FurtherConsiderations (by L. Wirth)
@@ A.1 Auxiliary Lemmas
is a random variable on the discrete measurable space $(\{\pm 1\}^{m},\mathcal{P}(\{\pm 1\}^{m}))$ with
@@
Furthermore, $V_{h}$ is an average of independent random variables, each of which takes values in [1,1][-1,1].
@@
where, for any $i\in[m]$, the random variable [^14] Note that the map $V_{h}^{(i)}$ is measurable, as it is defined on the discrete measurable space $(\{\pm 1\}^{m},\mathcal{P}(\{\pm 1\}^{m}))$. $V_{h}^{(i)}$ is given by
@@
$\mathbb{P}(V_{h}^{(i)}=0)$ $=\begin{cases}1&\text{if }\Delta_{h}^{(i)}=0,\\ 0&\text{otherwise},\end{cases}$
$\mathbb{P}(V_{h}^{(i)}=\ell)$ $=\begin{cases}\frac{1}{2}&\text{if }\Delta_{h}^{(i)}\neq 0,\\ 0&\text{otherwise},\end{cases}\quad\quad(\ell\in\{\pm 1\}),$
@@
The following result is a refined version of Shalev-Shwartz and Ben-David [28, Lemma A.4].
@@
If $\beta\geq e$, then applying [28, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=\beta$) yields
@@
Otherwise, applying [28, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=e$) yields
@@
Let $(\Omega_{1},\Sigma_{1})$$(\Omega_{2},\Sigma_{2})$ be measurable spaces and let $A_{1}\subseteq\Omega_{1}$, $A_{2}\subseteq\Omega_{2}$. Then we have $(\Sigma_{1}\otimes\Sigma_{2})\cap(A_{1}\times A_{2})=(\Sigma_{1}\cap A_{1})\otimes(\Sigma_{2}\cap A_{2})$.
@@
$=\;$ $\{(E_{1}\cap A_{1})\times(E_{2}\cap A_{2})\mid E_{1}\in\Sigma_{1},E_{2}\in\Sigma_{2}\}$
$=\;$ $\{E^{\prime}_{1}\times E^{\prime}_{2}\mid E^{\prime}_{1}\in(\Sigma_{1}\cap A_{1}),E^{\prime}_{2}\in(\Sigma_{2}\cap A_{2})\},$
@@ A.2 Well-Behaved Hypothesis Spaces
$\underline{z}$ $\mapsto\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|$
@@
is $\Sigma_{\mathcal{Z}}^{2m}$–measurable. In general, we however need to take suprema over uncountable families of these measurable functions. In this situation, verifying the measurability of the maps $U$ and $V$ is more complicated or even impossible (see Example A.13). However, in this section we aim at identifying further sufficient conditions that might be helpful in the uncountable setting as well.
We now introduce the concept of a universally separable hypothesis space. Dudley [11, page 902] was the first to introduce this property “as a way of avoiding measurability difficulties” (Pollard [26, page 38]).
@@
We next prove that universal separability ensures measurability of the maps $U$ and $V$ from Definition 3.2. Similar statements in a more general setting can be found in the literature (cf. e.g. Ben-David, Mansour and Benedek [3, Lemma 5.2] and Pollard [26, Chapter II, Problem 3]). Guided by the proof of [3, Lemma 5.2] and by arguments suggested by D. Pollard [^15] The second author thanks D. Pollard for these suggestions via private communication., we present here a detailed proof adapted to the maps $U$ and $V$.
@@
$\Leftrightarrow\,$ $\ h(x)\neq h_{k-1}(x)\text{ and }h(x)=h_{\ell}(x)\text{ for any }\ell\geq k$
$\Leftrightarrow\,$ $\ (x,y)\in(\Gamma(h)\triangle\Gamma(h_{k-1}))\text{ and}$
@@
$\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h_{n})-\operatorname{er}_{\mathbb{D}}(h)\big|=\big|\mathbb{D}(\Gamma(h))-\mathbb{D}(\Gamma(h_{n}))\big|\leq\mathbb{D}(\Gamma(h)\triangle\Gamma(h_{n})),$
@@
The universal separability condition applies to many standard examples of hypothesis spaces (cf. Pollard [26, Chapter II, Problems 4, 5 and 7]). However, universal separability is no necessary condition for ensuring well-behavedness. In fact, Example A.15 presents a quite simple hypothesis space that is not universally separable, but well-behaved with respect to certain sets of distributions.
Further considerations on the measurability of suprema over uncountable families of measurable functions can be found in Dudley [11], [12, Chapter 10], Gaenssler [13, Chapter 4] and Pollard [26, Appendix C]. For instance, [26, Appendix C] introduces the concept of permissible function classes, which Haussler [15, § 9.2] and Lee [19, § 2.3] also take up, e.g. to derive similar bounds as in Theorem 3.6. Recall that the arguments in [26, Appendix C], originally developed for permissible function classes, form the core for our proof of Theorem 4.7. Most hypothesis spaces arising in applications can in fact be shown to be permissible ([15, page 137]). The considerations in [26, Appendix C] further imply that permissibility constitutes a sufficient condition for well-behavedness. A permissible hypothesis space must be indexed by an analytic subset of a compact metric space, whereas well-behavedness reflects the minimal measure-theoretic requirements of the Fundamental Theorem of Statistical Learning, and is therefore maximally general. In theory, there could be non-permissible hypothesis spaces for which the more general framework of well-behavedness is nonetheless applicable.
@@
[…]
    $\text{and }\mathbb{D}^{m}(C)\geq 1-\delta.$
@@
[…]
@@
    for the finite set $C=\{\underline{z}\in C_{\mathbb{D}}^{m}\mid U(\underline{z})\leq\varepsilon(m,\delta)\}\in\Sigma_{\mathcal{Z}}$ and any $m\geq m_{0}$, where m_0=m_0(ε,δ)=⌈max{m_0^(1), m_0^(2),m_0^(3)}⌉.
@@
    $\subseteq\;$ $\left\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\varepsilon\right\}\!.$